OpenStudy (blake57roger):

What is the area of this circle? (Express your answer using π.) circle is 12 cm

OpenStudy (gorv):

pi*r*r radius =12 cm area=pi*12*12=144 *pi @blake57roger

OpenStudy (blake57roger):

thanks